OpenStudy (anonymous):

Consider the graph of a quadratic function shown to the right. A. Write a quadratic function that has x^2 -coefficient equal to 1 and has the x-intercepts indicated. B. Write a quadratic equation in the variable x, with X^2 -coefficienent equal to 1, that has solutions of -6 and -2 c. How many different quadratic functions can have x intercepts of -6 and -2? D. How many different quadratic equations in one variable can have solutions of -6, and -2?

OpenStudy (anonymous):

@Agent47

OpenStudy (anonymous):

The graph is a u shape and crosses on the x-axes at both -2, and -6.

OpenStudy (anonymous):

@Agent47

OpenStudy (agent47):

can u tell me at what y-value is its lowest point?

OpenStudy (anonymous):

-4 is it's lowest point on the y axis

OpenStudy (agent47):

brb gonna make some tea..

OpenStudy (anonymous):

ok

OpenStudy (agent47):

A. Write a quadratic function that has x^2 - coefficient equal to 1 and has the x-intercepts indicate: y=x^2+bx+c 0 Intercepts are -6 and -4, so: 0=(-6)^2+b(-6)+c 0=(-2)^2+b(-2)+c -36=-6b+c -4=-2b+c

OpenStudy (anonymous):

ok so that is going to be my A part for my formula?

OpenStudy (agent47):

A: y=x^2+8x+14

OpenStudy (agent47):

wait let me check it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=x^2+8x+12

OpenStudy (anonymous):

ok im back my computer knocked me off

OpenStudy (agent47):

sorry i froze when i was typing that 14 shld be 12

OpenStudy (agent47):

yeah Marina's right

OpenStudy (anonymous):

so it's x^2+8x+12?

OpenStudy (anonymous):

For a part

OpenStudy (agent47):

yea for A

OpenStudy (anonymous):

ok

OpenStudy (agent47):

B is the same thing

OpenStudy (anonymous):

same answer?

OpenStudy (agent47):

yea

OpenStudy (anonymous):

for the next one it says ___different quadratic functions or infintiy

OpenStudy (agent47):

c. How many different quadratic functions can have x intercepts of -6 and -2? infintely many

OpenStudy (agent47):

infinity

OpenStudy (anonymous):

Is D going to be infinity also?

OpenStudy (agent47):

D. How many different quadratic equations in one variable can have solutions of -6, and -2? yea it shld be infinity as well

OpenStudy (anonymous):

Great!!! Thanks alot :)

OpenStudy (agent47):

np gnite guys

OpenStudy (anonymous):

gnite